1. A box contains 3 red,10 white, 5 blue and 2 green balls identical in mass and shape. Find the ratio of the number of green ball to the total numbers of balls.



2. The number of goals scored by a famous footballer per match in a tournament are:


0,2,1 0,1,3,2,1,1,1,4,1,1,2,1 2,0,0,1,1,1,1,0,2,0,0,2,1,1,3


From this data, find the ratio of the number of matches where the footballer scored more than two goals to the total number of matches.


1
Expert's answer
2022-04-19T01:38:36-0400

"1:\\\frac{2}{3+10+5+2}=0.1\\2:\\0,2,\boldsymbol{1}\boldsymbol{0},1,\boldsymbol{3},2,1,1,1,\boldsymbol{4},1,1,2,\boldsymbol{1}\boldsymbol{2},0,0,1,1,1,1,0,2,0,0,2,1,1,\boldsymbol{3}\\\frac{5}{28}=0.178571"